OpenStudy (justme):

has anyone done college algebra 2?

OpenStudy (anonymous):

no haha im a lithin

OpenStudy (amistre64):

yes; maybe..no matrix stuff tho

OpenStudy (justme):

how do i solve this log 125 base x =3

OpenStudy (anonymous):

if you rewrite this in exponential form you have x^3 = 125, now either which number times itself three times is 125 or what is the third root of 125.

OpenStudy (justme):

5

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

that it!

OpenStudy (anonymous):

sometimes it's easier to solve a log problem if you rewrite it in exponential form.

OpenStudy (justme):

so on log problems, u make it equal. I got to the part 5^3 = x^3. thank u for your help.